# The Energy Sharing Timescale in an Analytic Framework for Common Envelope Hydrodynamics

### v1

@@ 1 Introduction
Extensive efforts have been undertaken to attempt to map pre-CE systems to their post-CE products (e.g., van den Heuvel, 1976; Davis et al., 2010; Kruckow et al., 2016; Vigna-Gómez et al., 2020; Marchant et al., 2021). Yet, CE evolution is complex, with key physical processes spanning a large range of length and time scales, requiring approximations in any mapping. The standard treatment is a decades-old energy formalism (see Section 2) which parameterizes a simple energy criterion for envelope ejection via an efficiency term (van den Heuvel, 1976; Webbink, 1984; Livio & Soker, 1988; de Kool, 1990; Iben & Livio, 1993). It is well known that this prescription for envelope ejection is incomplete, if not problematic (see, e.g., Soker, 1992, 2017; Clayton et al., 2017; Glanz & Perets, 2018; Ivanova et al., 2020), folding many physical processes into a few parameters when perhaps additional criteria should be included. In addition, there are known systems, such as close double white dwarfs and stellar-mass binary black holes, that the CE energy formalism cannot reproduce, though the use of angular momentum as an alternative criterion has brought some relief (Nelemans et al., 2000; Hirai & Mandel, 2022; Di Stefano et al., 2023).
@@ 2 The Energy Formalism
We begin by defining the standard CE energy formalism (van den Heuvel, 1976; Webbink, 1984; Livio & Soker, 1988; de Kool, 1990; Iben & Livio, 1993). The result of a CE phase in this model depends on the deposition of orbital energy into the binding energy of the envelope: if the energy change is at least equal to the binding energy, the envelope ejection is said to be successful. The orbital energy deposited due to orbital tightening is defined as
$\Delta E_{\mathrm{orb}}(r)=\frac{GM_{\mathrm{1}}M_{\mathrm{2}}}{2a_{\mathrm{init}}}-\frac{GM_{\mathrm{enc}}M_{\mathrm{2}}}{2r},$ (1)
@@
It is understood that not all of the orbital energy will contribute to the ejection of the envelope (Iaconi et al., 2018; Iaconi & De Marco, 2019), and it is likely that other sources of energy, e.g. recombination energy, may contribute (Iben & Livio, 1993; Ivanova et al., 2013). Therefore, this quantity is modified by an efficiency term $\alpha$ that functions as a “catch-all” for energetic contributions and deficits during a CE phase: if considering only orbital energy deposition, then $\alpha\leq 1$, but if folding in contributions from other energy sources, $\alpha$ can increase the available energy budget by up to a factor of $\approx 5$ (Sand et al., 2020). It is this modified quantity that must be sufficient to unbind the envelope; this defines the so-called $\alpha$-formalism:
@@
in which $E_{\mathrm{bind}}$ is the total binding energy of the envelope. For simplicity, we adopt $\alpha=1$ and do not include additional energy source terms nor the specific internal energy of the gas (though, see the discussion of Ivanova et al., 2015; Soker, 2017; Grichener et al., 2018; Soker et al., 2018; Reichardt et al., 2020), defining the total binding energy of the envelope as equivalent to the gravitational binding energy
$E_{\mathrm{grav}}=\int_{M_{\mathrm{core}}}^{M_{1}}-\frac{GM_{\mathrm{enc}}}{r}dm\equiv\frac{GM_{1}M_{\mathrm{env}}}{\lambda R},$ (3)
@@ 3.1 Energy: The Ejection Parameter, $\xi$
This adjustment to $E_{\mathrm{grav,SE}}$ is distinct in that it does not require the entire envelope to be ejected, but only the material exterior to $r-R_{\mathrm{a}}$, because this will be enough to effectively mitigate the drag force experienced by the secondary and halt its infall.
@@
In the standard energy formalism, the energy criterion for envelope ejection at and above a given position is met where $\alpha\geq 1$; analogously, the energy criterion for envelope ejection sufficient to end inspiral (hereafter, sufficient ejection) at a given position is met where $\xi\geq 1$.
@@ 3.2 Time: The Dynamical Parameter, $\beta$
We argue that $\xi$ alone cannot predict the outcome of a CE interaction. Because energy transport does not occur instantaneously throughout the envelope, assessing the degree to which energy is shared can improve our model. Shocks launched by gravitational interaction between the secondary and the surrounding gas are the mechanism by which energy is shared and orbital decay forces are generated. Of interest here are how the orbital decay rate compares to the rate that shocks are able to diffuse through the envelope, redistributing energy and angular momentum.
@@
in which the maximum velocity for the shock induced by the secondary ($v_{\mathrm{Kep}}$) and the sound speed of the gas at $r$ ($c_{\mathrm{s}}$) provide lower and upper limits. This is a measure of the traversal time of the heat source in the envelope.
@@ 4.1 Stellar Profiles
In general, we base our model of orbital infall, binding energy, and energy redistribution on the properties of the stellar envelope of the primary star. In order to model representative envelope structures, we use a replicated setup of MESA v7503 (Paxton et al., 2011, 2013, 2015) containing the inlists and settings of the MIST package (Choi et al., 2016; Dotter, 2016) for primaries with $Z=Z_{\odot}$ (appropriate for modeling galactic CE events) and, due to limitations in MIST at low metallicity and high mass, MESA v10398 (Paxton et al., 2011, 2013, 2015, 2018) for primaries with $Z=1/50Z_{\odot}$ (appropriate for modeling Population III binaries, i.e. progenitors of binary gravitational wave sources).
@@ 5 Discussion
It has long been the goal of the CE community to develop a prescription that can approximate the results of the best state-of-the-art 3D models with the fewest inputs and at the smallest computational expense possible. Following previous efforts to improve upon the widely-utilized standard energy formalism (e.g., de Kool, 1990; Ivanova & Chaichenets, 2011; Hirai & Mandel, 2022; Di Stefano et al., 2023), this framework is the first to implement a component predictive of CE hydrodynamic wave speeds into an analytical model.
@@ 5.1 Additional Considerations
Further, we suggest that additional consideration will need to be included to best describe the “stopping” criteria of events. For example, the CE systems detailed in this work generally have primary cores massive enough to avoid disruption by the secondary during inspiral, but this is not always the case. Some treatment for the tidal disruption of either the primary or the secondary should be included to enhance the framework’s predictive capabilities (Ivanova et al., 2002). Secondly, when a CE achieves the ejection criteria of $\xi>1$ and $\beta>1$, it may leave behind some of the H-rich envelope interior to the current separation. This is not necessarily problematic: if the remaining envelope contracts onto the core any ongoing interaction will “shut off.” On the other hand, if the remaining envelope expands as the overlying layers are removed it would drive continued interaction (Soberman et al., 1997; Ivanova, 2011; Ivanova & Nandez, 2016).
@@ 5.2 Summary and Conclusion
We thank D. Lin, M. Gallegos-Garcia, J. Law-Smith, A. Loeb, S. Toonen, and R. Yarza for helpful discussions. R.W.E. acknowledges the support of the University of California President’s Dissertation-Year Fellowship, the Heising-Simons Foundation, and the Vera Rubin Presidential Chair for Diversity at UCSC. M.M. gratefully acknowledges support from a Clay Fellowship at the Smithsonian Astrophysical Observatory. E.R.-R. acknowledges support from the Heising-Simons Foundation and NSF (AST-2307710, AST-2206243, AST-1911206, and AST-1852393). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the NSF. This work was supported in part by NASA ATP Grant 80NSSC22K0722 to the University of North Carolina at Chapel Hill. The 3D hydrodynamics software used in this work was developed in part by the DOE NNSA- and DOE Office of Science-supported Flash Center for Computational Science at the University of Chicago and the University of Rochester. We acknowledge use of the computational resources of the SCIENCE HPC Center at the University of Copenhagen, with the support of the DARK Cosmology Centre at the Niels Bohr Institute.

### v2

@@ 1 Introduction
Extensive efforts have been undertaken to attempt to map pre-CE systems to their post-CE products (e.g., van den Heuvel, 1976; Davis et al., 2010; Kruckow et al., 2016; Vigna-Gómez et al., 2020; Marchant et al., 2021). Yet, CE evolution is complex, with key physical processes spanning a large range of length and time scales, requiring approximations in any mapping. The standard treatment is a decades-old energy formalism (see Section 2) which parameterizes a simple energy criterion for envelope ejection via an efficiency term (van den Heuvel, 1976; Webbink, 1984; Livio & Soker, 1988; de Kool, 1990; Iben & Livio, 1993). It is well known that this prescription for envelope ejection is incomplete, if not problematic (see, e.g., Soker, 1992, 2017; Clayton et al., 2017; Glanz & Perets, 2018; Ivanova et al., 2020), folding many physical processes into a few parameters when perhaps additional criteria should be included. In addition, there are known systems, such as close double white dwarfs and stellar-mass binary black holes, that the CE energy formalism cannot reproduce, though the use of convection or angular momentum as an alternative criterion has brought some relief (Nelemans et al., 2000; Wilson & Nordhaus, 2020; Hirai & Mandel, 2022; Wilson & Nordhaus, 2022; Di Stefano et al., 2023).
@@ 2 The Energy Formalism
We begin by defining the standard CE energy formalism (van den Heuvel, 1976; Webbink, 1984; Livio & Soker, 1988; de Kool, 1990; Iben & Livio, 1993). Essentially, this formalism depends on the key assumption of energy conservation, which requires that the timescale of a CE event be much shorter than the thermal timescale of the envelope such that additional energy sources and sinks can be ignored. However, in cases in which the inspiral timescale approaches or exceeds the thermal timescale, energy conservation becomes less reliable as nuclear energy generation starts dominating the energy budget. Likewise, cases in which the inspiral timescale exceeds the envelope’s convection timescale may be dominated by radiation losses (Wilson & Nordhaus, 2019, 2020, 2022). The result of a CE phase in this model depends on the deposition of orbital energy into the binding energy of the envelope: if the energy change is at least equal to the binding energy, the envelope ejection is said to be successful. The orbital energy deposited due to orbital tightening is defined as
$\Delta E_{\mathrm{orb}}(r)=\frac{GM_{\mathrm{enc}}M_{\mathrm{2}}}{2r}-\frac{GM_{\mathrm{1}}M_{\mathrm{2}}}{2a_{\mathrm{init}}},$ (1)
@@
It is understood that not all of the orbital energy will contribute to the ejection of the envelope (Iaconi et al., 2018; Iaconi & De Marco, 2019), and it is likely that other sources of energy may contribute (Iben & Livio, 1993; Ivanova et al., 2013), e.g. recombination energy (though this has been rigorously disputed, see Sabach et al., 2017; Grichener et al., 2018). Therefore, this quantity is modified by an efficiency term $\alpha$ that functions as a “catch-all” for energetic contributions and deficits during a CE phase: if considering only orbital energy deposition, then $\alpha\leq 1$, but if folding in contributions from other energy sources, $\alpha$ can increase the available energy budget by up to a factor of $\approx 5$ (Sand et al., 2020). It is this modified quantity that must be sufficient to unbind the envelope; this defines the so-called $\alpha$-formalism:
@@
in which $E_{\mathrm{bind}}$ is the total binding energy of the envelope. For simplicity, we adopt $\alpha=1$ and do not include additional energy source terms nor the specific internal energy of the gas (though, see the discussion of Ivanova et al., 2015; Soker, 2017; Sabach et al., 2017; Grichener et al., 2018; Soker et al., 2018; Reichardt et al., 2020), defining the total binding energy of the envelope as equivalent to the gravitational binding energy
[…]
@@ 3.1 Energy: The Ejection Parameter, $\xi$
This adjustment to $E_{\mathrm{grav,SE}}$ is distinct in that it does not require the entire envelope to be ejected, but only the material exterior to $r-R_{\mathrm{a}}$, because this will be enough to effectively mitigate the drag force experienced by the secondary and halt its infall in envelope regions that contract in response to mass loss, or significantly slow its infall due to reduced drag in envelope regions that expand in response to mass loss.
@@
In the standard energy formalism, the energy criterion for envelope ejection at and above a given position is met where $\alpha\geq 1$; analogously, the energy criterion for envelope ejection sufficient to end inspiral (hereafter, sufficient ejection) at a given position is met where $\xi\geq 1$. We note that the assumption of energy conservation native to the standard CE energy formalism is maintained here, therefore nuclear energy generation and radiative losses are not considered.
@@ 3.2 Time: The Dynamical Parameter, $\beta$
We argue that $\xi$ alone cannot predict the outcome of a CE interaction. Because energy transport does not occur instantaneously throughout the envelope, assessing the degree to which energy is shared can improve our model. Shocks launched by gravitational interaction between the secondary and the surrounding gas are the mechanism by which energy is shared and orbital decay forces are generated (for a discussion of the analogous role of pressure waves in very low mass ratio systems, see Soker, 1993). Of interest here are how the orbital decay rate compares to the rate that shocks are able to diffuse through the envelope, redistributing energy and angular momentum.
@@
in which the maximum velocity for the shock induced by the secondary ($v_{\mathrm{Kep}}$) and the sound speed of the gas at $r$ ($c_{\mathrm{s}}$) provide lower and upper limits [^1] These limits apply to lateral energy sharing; however the convective velocity may in some cases be comparable to these wave speeds, leading to energy losses through radial transport to the photosphere. This would have no impact on $\beta$, but would effectively reduce $\xi$.. This is a measure of the traversal time of the heat source in the envelope.
@@ 4.1 Stellar Profiles
In general, we base our model of orbital infall, binding energy, and energy redistribution on the properties of the stellar envelope of the primary star. In order to model representative envelope structures, we use a replicated setup of MESA v7503 (Paxton et al., 2011, 2013, 2015) containing the inlists and settings of the MIST package (Choi et al., 2016; Dotter, 2016) for primaries with $Z=Z_{\odot}$ (appropriate for modeling galactic CE events) and, due to limitations in MIST at low metallicity and high mass, MESA v10398 (Paxton et al., 2011, 2013, 2015, 2018) for primaries with $Z=1/50Z_{\odot}$ (appropriate for modeling Population III binaries, i.e. progenitors of binary gravitational wave sources) [^2] The inlists, source files, outputs, and analysis code necessary to reproduce the models and figures in this work are available on Zenodo: https://doi.org/10.5281/zenodo.14284113 (catalog doi:10.5281/zenodo.14284113).
@@ 5 Discussion
It has long been the goal of the CE community to develop a prescription that can approximate the results of the best state-of-the-art 3D models with the fewest inputs and at the smallest computational expense possible. Following previous efforts to improve upon the widely-utilized standard energy formalism (e.g., de Kool, 1990; Ivanova & Chaichenets, 2011; Wilson & Nordhaus, 2019; Hirai & Mandel, 2022; Wilson & Nordhaus, 2022; Di Stefano et al., 2023; Noughani et al., 2024), this framework is the first to implement a component predictive of CE hydrodynamic wave speeds into an analytical model.
@@ 5.1 Additional Considerations
At higher $q$ values, the companion’s gravitational reach may be better described by the Roche radius (Eggleton, 1983) than $R_{\mathrm{a}}$. Such considerations are relevant in higher $q$ binaries, as well as late stages of CE when the secondary nears the primary’s core, as the assumptions behind the drag formalism break down when $R_{\mathrm{a}}$ penetrates the core. Therefore, the calculation for $\xi$ would benefit from a more sophisticated treatment for such cases. In fact, CE wind tunnel simulations (MacLeod & Ramirez-Ruiz, 2015a; MacLeod et al., 2017; De et al., 2020) show that the presence of a steep density gradient sharply redirects shock-heated material outward as waves couple more easily with less dense material, suggesting that the core may be preserved and no part of its binding energy need be included in $E_{\mathrm{grav,SE}}$, giving higher $\xi$ values in this region.
Further, we suggest that additional consideration will need to be included to best describe the “stopping” criteria of events. For example, the CE systems detailed in this work generally have primary cores massive enough to avoid disruption by the secondary during inspiral, but this is not always the case. Some treatment for the tidal disruption of either the primary or the secondary should be included to enhance the framework’s predictive capabilities (Ivanova et al., 2002). Secondly, when a CE achieves the ejection criteria of $\xi>1$ and $\beta>1$, it may leave behind some of the H-rich envelope interior to the current separation. This is not necessarily problematic: if the remaining envelope contracts onto the core any ongoing interaction will “shut off.” On the other hand, if the remaining envelope expands as the overlying layers are removed it would drive continued interaction (Soberman et al., 1997; Ivanova, 2011; Ivanova & Nandez, 2016; Vigna-Gómez et al., 2020).
@@ 5.2 Summary and Conclusion
We thank D. Lin, M. Gallegos-Garcia, N. Soker, J. Law-Smith, A. Loeb, S. Toonen, and R. Yarza for helpful discussions. We also gratefully acknowledge the anonymous referee for their insightful feedback. R.W.E. acknowledges the support of the University of California President’s Dissertation-Year Fellowship, the Heising-Simons Foundation, and the Vera Rubin Presidential Chair for Diversity at UCSC. M.M. gratefully acknowledges support from a Clay Fellowship at the Smithsonian Astrophysical Observatory. E.R.-R. acknowledges support from the Heising-Simons Foundation and NSF (AST-2307710, AST-2206243, AST-1911206, and AST-1852393). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the NSF. This work was supported in part by NASA ATP Grant 80NSSC22K0722 to the University of North Carolina at Chapel Hill. The 3D hydrodynamics software used in this work was developed in part by the DOE NNSA- and DOE Office of Science-supported Flash Center for Computational Science at the University of Chicago and the University of Rochester. We acknowledge use of the computational resources of the SCIENCE HPC Center at the University of Copenhagen, with the support of the DARK Cosmology Centre at the Niels Bohr Institute.